# Pneumatic Multi-mode Silicone Actuator with Pressure, Vibration, and Cold Thermal Feedback

### v1

@@ VII-A Stimuli
VII-B Participants and Experimental Design
These experiments had the same participants as User Study 1. All participants gave their informed written consent to participate in this study. All experiments were approved by the Institutional Review Board (IRB) from the author’s institution. A within-subject design was employed for the study. For each participant, the exploration on the surfaces using the two systems was repeated twice, yielding 8 explorations per participant. They were also asked to push the virtual 3D button using the same two actuators. Pushing the 3D button using both actuators was repeated five times, yielding 10 explorations per participant.
VII-C Experimental Procedure
Before starting the experiment, the participants were instructed about the procedure. We confirmed that participants correctly understood the procedure by asking them to repeat the instructions. This experiment was divided into training and main sessions.
In the training session, participants were introduced to the two types of actuators and the virtual environments. During the main session, participants were sound-blocked while they explored the virtual 3D surfaces and the virtual button using both actuators. The order of presentation was randomized. After exploring the surfaces and the button, participants were asked to rate the overall feedback fidelity by providing scores to the operator in response to the following questions: a) How realistic was the feedback? (realism), b) Was the feedback unnatural? (reverse scale), and c) How responsive was the feedback to actions in the VR environment? Ratings were given on a scale from 0 to 100 (0 meaning ”Not at all” and 100 meaning ”Very much”). The entire experiment took approximately thirty minutes to complete for each participant.
VII-D Result and Discussion
@@
For the texture scenarios, no significant realism difference was observed. We speculate that although the pneumatic actuator uses a less sophisticated rendering algorithm compared to the state-of-the-art algorithm for the voice-coil, it still preserves the comparable realism due to the additional cue provided by static pressure for contact. However, further investigation is required to pinpoint the causes. Additionally, we noted that the voice coil actuator performed slightly better on the fine rough surface (Figure 14(b)), likely due to its superior high-frequency rendering capability.
@@
VIII User Study 3: Overall Realism Assessment
@@
VIII-A VR Environments
To demonstrate the effectiveness of the proposed multi-mode actuator, we implemented two VR application scenarios. The first scenario involves touching a frozen meat surface, allowing participants to experience simultaneous sensations of contact pressure, cold thermal feedback, and surface stiffness. The second scenario features interaction with an abrasive icy surface, enabling the perception of contact pressure, cold thermal feedback, and surface texture. The virtual environments were developed using the Unity game engine (version 2020.3.43f1). The user’s finger was tracked using an optical tracker (OptiTrack Trio, V120, NaturalPoint, Inc., Corvallis, OR, USA) [44], and the tracked finger was visually represented in the scenes.
VIII-A1 Touching Frozen Meat Surface: simultaneous cold, pressure, and stiffness feedback
This scenario enables participants to perceive contact pressure, cold thermal feedback, and surface stiffness while interacting with a frozen meat surface. This section outlines the control schemes and rendering algorithms used to deliver these three simultaneous feedback modalities.
Figure 15: System for rendering frozen meat surface.
Figure 15 illustrates the rendering hardware and the virtual scene. When the user makes contact with the surface, both contact pressure and thermal feedback are activated, simulating the sensation of touching a cold surface. Simultaneously, dynamic pressure is applied in response to the user pressing on the frozen meat, creating a realistic perception of stiffness.
@@ VIII-A1 Touching Frozen Meat Surface: simultaneous cold, pressure, and stiffness feedback
2:  while isTouched($P$,$MeatMesh$do
@@
5:     $k\leftarrow$ getTargetStiffness($MeatMesh$) {Retrieve target stiffness}
@@
7:     $T\leftarrow$ getTargetTemperature($MeatMesh$) {Retrieve target temperature}
@@
10:
11:  Finish
@@
VIII-A2 Touching Abrasive Icy Surface: Simultaneous Cold Thermal, Contact Pressure, and Texture Feedback
This scenario is designed to allow the user to perceive contact pressure, cold thermal feedback, and the texture of surface roughness while stroking across the surface. Figure 16 displays the setup and the VR scene. For texture rendering, the grid-based algorithm introduced VII is utilized.
Figure 16: System for rendering abrasive icy surface.
@@ VIII-A2 Touching Abrasive Icy Surface: Simultaneous Cold Thermal, Contact Pressure, and Texture Feedback
2:  while isTouched($P$, $IcyMesh$do
@@
4:     RenderPressure($cp$) {Render constant pressure for contact feeling}
5:     $G\leftarrow$ getTargetTextureGrid($IcyMesh$) {Retrieve the texture grid structure of the target surface}
6:     if isPassingGrid($P$, $G$then
@@
10:     $T\leftarrow$ getTargetTemperature($IcyMesh$) {Retrieve target temperature}
@@
13:
14:  Finish
The rendering algorithm in Algorithm 2 closely resembles that of Algorithm 1, with the primary distinction being in the vibration generation for texture feedback (lines 6-9). The texture rendering part first retrieves the grid structure from the contacting mesh and checks if the user’s hand is passing over one of the grids. If so, it generates short vibration bursts to simulate the acceleration associated with surface roughness. The faster the user’s movement, the stronger the vibration experienced.
VIII-B Participants
A subset of ten participants (8 males, 2 females; average age 29) from the previous experiment participated in this user study. Seven had no experience with haptic rendering, and three had general knowledge. None reported conditions affecting hand sensations. All participants gave their informed written consent to participate in this study. All experiments were approved by the Institutional Review Board (IRB) from the author’s institution.
VIII-C Experimental Design
Four haptic feedback conditions were the subject of a within-subject experiment: a) No feedback (NF), b) Tactile feedback (TF), c) Cold thermal feedback (CTF), and d) Simultaneous thermal-tactile feedback (STF). No feedback refers to a condition in which there was neither cold thermal feedback nor tactile haptic feedback. Tactile feedback presented stiffness plus contact pressure feedback for the cold meat surface, and texture plus contact pressure feedback for the abrasive icy surface, without activating the cold thermal feedback. Cold thermal feedback presented just cold thermal feedback, with no tactile feedback. Finally, Simultaneous thermal-tactile feedback offered both tactile and cold thermal feedback simultaneously.
@@ VIII-C Experimental Design
    Localization – The precise location of the haptic feedback was identifiable within the VR environment.
@@
    Reaction – The system synchronized effectively with user interactions in the VR setting.
@@
After each condition, participants rated each question on a continuous scale from 0 to 100, with ”Not at all” and ”Very much” as the endpoints for each feedback condition.
VIII-D Experimental Procedure
@@
Figure 17: Mean scores and standard error of all the measures in different feedback conditions for frozen meat surface.
Figure 18: Mean scores and standard error of all the measures in different feedback conditions for abrasive icy surface
VIII-E Result and Discussion
Figures 17 and 18 illustrate the average scores and standard errors for the frozen meat surface and abrasive icy surface scenarios respectively.
@@ VIII-E Result and Discussion
A one-way ANOVA with repeated measurements was used to analyze the collected data. Post hoc comparison using the Tukey test revealed a significant difference (p $<$ 0.01) between Simultaneous cold thermal-tactile feedback and No feedback for the first, second, and fifth measures in both scenarios. Furthermore, a significant difference was observed between Tactile feedback and Simultaneous thermal-tactile feedback for the second and third measures (p $<$ 0.05) on the meat surface scenario, while for the first, second, and fifth measures on the abrasive icy surface scenario. However, no significant difference was found between Tactile feedback and Cold thermal feedback across any measures in either scenario. Moreover, Users rated the third measure in the No feedback condition as zero due to the lack of tactility in both scenarios.
@@
The results indicate that integrating tactile and cold thermal cues can significantly enhance the virtual experience. In both VR environments, Simultaneous thermal-tactile feedback was the preferred choice across all measures. Conversely, no feedback was consistently rated the least favorably, while all other feedback conditions were preferred over this baseline. Overall, our multi-modal approach proves beneficial in increasing haptic realism in suitable interaction scenarios.
IX Conclusion
@@
This work introduces a silicone-based fingertip actuator designed to replicate realistic tactile sensations by delivering multiple types of haptic feedback. The actuator simultaneously provides thermal, pressure, and vibrotactile feedback through dual air chambers and lateral air nozzles. This design enables users to experience subtle temperature variations and tactile sensations (pressure and vibration) that closely mimic real-world interactions.
@@
The actuator is adaptable, accommodating various finger sizes and featuring adjustable air chambers to control feedback intensity. Additionally, pressure modulation based on depth enhances realism during interactions with virtual surfaces, as demonstrated in successful tests on cold meat surfaces.
@@
This approach serves as an effective alternative to conventional wearable haptic devices, which often struggle to provide comprehensive multi-modal feedback using a single actuator. Its potential applications span education, safety training, medical training, and rehabilitation, enriching the realism of virtual training experiences. The actuator enhances haptic technology with its flexible and efficient design, ultimately improving user experiences and contributing to advancements in virtual and augmented reality systems.

### v2

@@ VII-B1 Stimuli
VII-B2 Participants and Experimental Design
This experiment had the same participants as User Study 1. A within-subjects design was employed for the study. The exploration on the surfaces using the two systems was repeated twice for each participant, yielding 8 explorations per participant. They were also asked to push the virtual 3D button using the same two actuators. Pushing the 3D button using both actuators was repeated five times, yielding 10 explorations per participant.
VII-B3 Experimental Procedure
Before starting the experiment, participants were instructed about the procedure. We confirmed that participants correctly understood the procedure by asking them to repeat the instructions. This experiment was divided into training and main sessions. In the training session, participants were introduced to the two types of actuators and the virtual environments.
During the main session, participants were sound-blocked while exploring the virtual 3D surfaces and the virtual button using both actuators. The order of presentation was randomized. After exploring the surfaces and the button, participants were asked to rate the overall feedback fidelity by providing scores to the operator in response to the following questions: a) How realistic was the feedback? (realism), b) Was the feedback unnatural? (reverse scale), and c) How responsive was the feedback to actions in the VR environment? Ratings were given on a scale from 0 to 100 (0 meaning “Not at all” and 100 meaning “Very much”). The entire experiment took around 30 minutes to complete for each participant.
VII-B4 Result and Discussion
@@
For the texture scenarios, no significant realism difference was observed. We speculate that, although the pneumatic actuator employs a less sophisticated rendering algorithm compared to the state-of-the-art algorithm for the voice coil, it still preserves comparable realism due to the additional cue provided by static pressure for contact. However, further investigation is required to pinpoint the causes. Additionally, we noted that the voice coil actuator performed slightly better on the fine rough surface (Figure 14(b)), likely due to its superior high-frequency rendering capability.
@@
VII-C User Study 3: Overall Realism Assessment
@@
VII-C1 VR Environments
To demonstrate the effectiveness of the proposed multi-mode actuator, we implemented two VR application scenarios. The first scenario involves touching a frozen meat surface, allowing participants to experience simultaneous sensations of contact pressure, cold thermal feedback, and surface stiffness. The second scenario features interaction with an abrasive icy surface, enabling the perception of contact pressure, cold thermal feedback, and surface texture. The virtual environments were developed using the Unity game engine (version 2020.3.43f1). The user’s finger was tracked using an optical tracker (OptiTrack Trio, V120, NaturalPoint, Inc., Corvallis, OR, USA) [OptiTrackV120], and the tracked finger was visually represented in the scenes.
Figure 15: System for rendering (a) frozen meat surface, and (b) abrasive icy surface.
@@ VII-C1 VR Environments
2:while isTouched($P$,MeatMeshMeatMesh) do
@@
5:  $k\leftarrow$ getTargetStiffness(MeatMeshMeatMesh) {Retrieve target stiffness}
@@
7:  $T\leftarrow$ getTargetTemperature(MeatMeshMeatMesh) {Retrieve target temperature}
@@
10:Finish
i) Touching Frozen Meat Surface (simultaneous cold, pressure, and stiffness feedback): This scenario enables participants to perceive contact pressure, cold thermal feedback, and surface stiffness while interacting with a frozen meat surface. This section outlines the control schemes and rendering algorithms used to deliver these three simultaneous feedback modalities.
Figure 15(a) illustrates the rendering hardware and the virtual scene. When the user makes contact with the surface, both contact pressure and thermal feedback are activated, simulating the sensation of touching a cold surface. Simultaneously, dynamic pressure is applied in response to the user pressing on the frozen meat, creating a realistic perception of stiffness.
@@
2:while isTouched($P$, IcyMeshIcyMesh) do
@@
4:  RenderPressure(cpcp) {Render constant pressure for contact feeling}
5:  $G\leftarrow$ getTargetTextureGrid(IcyMeshIcyMesh) {Retrieve the texture grid structure of the target surface}
6:  if isPassingGrid($P$, $G$) then
@@
10:  $T\leftarrow$ getTargetTemperature(IcyMeshIcyMesh) {Retrieve target temperature}
@@
13:Finish
ii) Touching Abrasive Icy Surface (Simultaneous Cold Thermal, Contact Pressure, and Texture Feedback): This scenario is designed to allow the user to perceive contact pressure, cold thermal feedback, and the texture of surface roughness while stroking across the surface. Figure 15(b) displays the setup and the VR scene. For texture rendering, the grid-based algorithm introduced VII-B is utilized.
The rendering algorithm in Algorithm 2 closely resembles that of Algorithm 1, with the primary distinction in vibration generation for texture feedback (lines 5–9). The texture rendering part first retrieves the grid structure from the contacting mesh and checks if the user’s hand is passing over one of the grids. If so, it generates short vibration bursts to simulate acceleration associated with surface roughness. The faster the user’s movement, the stronger the vibration experienced.
VII-C2 Participants
A subset of ten healthy participants (8 males, 2 females; average age 29) from the last experiment participated in this user study.
VII-C3 Experimental Design
We tested four haptic feedback conditions in a within-subject experiment: a) No feedback (NF), b) Tactile feedback (TF), c) Cold thermal feedback (CTF), and d) Simultaneous thermal-tactile feedback (STF). No feedback refers to a condition in which there was neither cold thermal feedback nor tactile haptic feedback. Tactile feedback presented stiffness plus contact pressure feedback for the cold meat surface, and texture plus contact pressure feedback for the abrasive icy surface, without activating the cold thermal feedback. Cold thermal feedback presented just cold thermal feedback, with no tactile feedback. Finally, Simultaneous thermal-tactile feedback offered both tactile and cold thermal feedback simultaneously.
@@ VII-C3 Experimental Design
    Localization – The precise location of haptic feedback was identifiable within the VR environment.
@@
    Reaction – The system was effectively synchronized with user interactions in the VR setting.
@@
After each condition, participants rated each question on a continuous scale from 0 to 100, with “Not at all” and “Very much” as endpoints for each condition.
Figure 16: Mean scores and standard error of all the measures in different feedback conditions for (a) frozen meat surface; (b)abrasive icy surface.
VII-C4 Experimental Procedure
@@
VII-C5 Result and Discussion
@@
Figures 16(a) and 16(b) illustrate the average scores and standard errors for the frozen meat surface and abrasive icy surface scenarios, respectively. A one-way ANOVA with repeated measurements was used to analyze the collected data. Post hoc comparison using the Tukey test revealed a significant difference ($p<0.01$) between Simultaneous cold thermal-tactile feedback and No feedback for the first, second, and fifth measures in both scenarios. Furthermore, a significant difference was observed between tactile feedback and Simultaneous thermal-tactile feedback for the second and third measures ($p<0.05$) in the meat surface scenario, while for the first, second, and fifth measures on the abrasive icy surface scenario. However, no significant differences were found between tactile feedback and cold thermal feedback in any measure in either scenario. Moreover, users rated the third measure in the No feedback condition as zero due to the lack of tactility in both scenarios.
@@
The results indicate that the integration of tactile and cold thermal cues can significantly enhance the virtual experience. In both VR environments, Simultaneous thermal-tactile feedback was the preferred choice across all measures. In contrast, no feedback was consistently rated less favorably, while all other feedback conditions were preferred over this baseline. Overall, our multi-modal approach proves beneficial in increasing haptic realism in suitable interaction scenarios.
VIII Conclusion and Future Work
@@
This work introduces a silicone-based fingertip actuator to replicate realistic tactile sensations by delivering multiple types of haptic feedback. The actuator simultaneously provides cold thermal, pressure, and vibrotactile feedback through dual air chambers and lateral air nozzles. This design enables users to experience subtle temperature variations and tactile sensations (pressure and vibration) that closely mimic real-world interactions.
@@
The actuator is adaptable, accommodating various finger sizes and features, with adjustable air chambers to control feedback intensity. In addition, modulation of the pressure based on depth enhances realism during interactions with virtual surfaces, as demonstrated in successful tests on cold meat surfaces.
@@
This approach serves as an effective alternative to conventional wearable haptic devices, which often struggle to provide comprehensive multi-modal feedback using a single actuator. Its potential applications span education, safety training, medical training, and rehabilitation, enriching the realism of virtual training experiences.
However, the actuator’s thermal rendering performed well in short-duration tests, but long-term stability could not be evaluated due to limitations in the air supply system. The vortex tube requires a continuous input pressure above 5.5 bar to maintain effective cooling, but the available source could only sustain this pressure for slightly more than 10 seconds. The cooling performance decreases as the pressure drops, limiting cold feedback to short periods. This issue is related to the experimental setup rather than the actuator design. Future work will address this by using a high-capacity or pressure-regulated air supply to enable stable and extended thermal operation.
As a future direction, a heating element can be integrated alongside the existing cooling function to enable bidirectional thermal feedback, allowing the system to produce both warm and cold sensations for more realistic temperature rendering. In addition, the current actuator can be extended into a glove-based system that provides haptic feedback to all five fingers using lightweight, pneumatically controlled modules capable of synchronized pressure, vibration, and thermal output.
[Uncaptioned image] Mohammad Shadman Hashem received the B.S. degree in Mechanical Engineering from the Chittagong University of Engineering and Technology (CUET), Chittagong, Bangladesh, in 2018, and the M.S. degree in Computer Science and Engineering from Kyung Hee University, South Korea, in 2022. He is currently pursuing a Ph.D. degree in the Department of Computer Science and Engineering at Kyung Hee University, South Korea. His research interests include the design of wearable tactile actuators and haptic rendering.
[Uncaptioned image] Ahsan Raza received the B.S. degree in Computer Engineering from the University of Engineering and Technology (UET) Taxila, Taxila, Pakistan, in 2015, and a Ph.D. degree in Computer Science and Engineering from Kyung Hee University, South Korea, in 2024. His research interests include mid-air haptic feedback, haptic actuator design, and psychophysics.
[Uncaptioned image] Sama E Shan received the B.S. degree in Mechanical Engineering from the Chittagong University of Engineering and Technology (CUET), Chittagong, Bangladesh, in 2019, and M.S. degree in Computer Science and Engineering from Kyung Hee University, South Korea, in 2024. His research interests include the design of wearable tactile actuators and haptic rendering.
[Uncaptioned image] Seokee Jeon (Member, IEEE) received the B.S. and Ph.D. degrees in Computer Science and Engineering from the Pohang University of Science and Technology (POSTECH), in 2003 and 2010, respectively. He was a Postdoctoral Research Associate with the Computer Vision Laboratory, ETH Zürich. In 2012, he joined the Department of Computer Engineering at Kyung Hee University as an Assistant Professor. His research interests include haptic rendering in an augmented reality environment, and the usability of augmented reality applications.